The spring-flex exercise system consists of a spring with one end fixed and a handle on the other end. The idea is that you exercise your muscles by stretching the spring from its natural length, which is 31 cm. If a 200 Newton force is required to keep the spring stretched to a length of 47 cm, how much work is required to stretch it from 46 cm to 70 cm?
1
Expert's answer
2020-07-29T11:28:59-0400

The spring-flex exercise system consists of a spring with one end fixed and a handle on the other end. The idea is that you exercise your muscles by stretching the spring from its natural length, which is 31 cm. If a 200 Newton force is required to keep the spring stretched to a length of 47 cm, how much work is required to stretch it from 46 cm to 70 cm?


Let's make the following denominations:

L1 = 31 cm (natural length)

L2 = 47 cm

F = 200 N

A1 = 2 J

L3 = 46 cm

L4 = 70 cm


We need to find the stiffness of the spring:

x1 = L2-L1 = 47 - 31 = 16 cm;

F = k*x1==> k = F/x1 = 200/16 = 25/2.


Let's find the work needed to stretch a spring from 31 cm to 46 cm:

x2 = L3-L1 = 46 - 31 = 15 cm

A2 = k*x2²/2 = 25/2 * 15² / 2 = 5625/4 J.


Let's find the work needed to stretch a spring from 31 cm to 70 cm:

x3 = L4-L1 = 70 - 31 = 39 cm

A3 = k*x3²/2 = 25/2 * 29² / 2 = 21025/4 J.


So, the work needed to stretch a spring from 60 cm to 74 cm is

A = A3 - A2 = 21025/4 - 5625/4 = 3850 J.